# Multi-Scale Data Assimilation in Turbulent Models

### v1

@@ V.2 Measured Scale Schemes
While unresolved scales have traditionally been treated as mere sources of model error [34, 35], in this work we perform a systematic investigation of how the choice of measured scales affects both statistical and instantaneous DA reconstruction, as well as cascading energy transfer. After establishing a synchronization threshold based on observation sparseness $\Delta t_{\textit{obs}}$, we fix it to $\tau_{15}$ (the turnover time of the dissipative scale), which so far has ensured the best assimilation performance, and vary the set of measured shells to assess its impact.
@@
Figure 7: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$ for different sets of measured shells; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(15)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
Figure 8: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$ for different sets of measured shells; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(15)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
@@
A central finding is that measuring two adjacent shells within the inertial range, such as shells 6 and 7 (Fig. 7), suffices to achieve full synchronization of the large-scale dynamics. Adding a third measurement (e.g., shells 6, 7, and 8) only assimilates an additional fast scale without significantly enhancing synchronization. In contrast, measuring a single shell (e.g., shell $6$) fails to synchronize the inertial range and disrupts the energy cascade. This disruption arises from the structure of the energy flux, which is mediated through triadic couplings involving terms such as $u_{n}^{*}u_{n+1}^{*}u_{n+2}$ and $u_{n-1}^{*}u_{n}^{*}u_{n+1}$. An inadequate reconstruction of these nonlinear interactions leads to bottlenecks, blocking the transfer of energy from the forced scales and causing its accumulation in the shells immediately preceding the measured one. Contiguity between measured shells is therefore critical: adjacent measurements fully constrain the local flux, ensuring seamless energy transfer across scales. When measurements are widely separated, synchronization gaps emerge, as illustrated in Fig. 8, where increasing separation in the 6–14, 4–14, and 2–14 configurations progressively deteriorates intermediate-scale synchronization. In the 2–14 case, a partial recovery is observed, attributed to the proximity of the (measured) $u_{2}$ shell to the forcing region (we recall that both the ground truth and the ensembles evolve with the same forcing), which improves local flux reconstruction. Although nonlocal correlations allow limited information transfer, the most effective strategy remains the measurement of two contiguous shells, simultaneously constraining both flux directions and ensuring full synchronization of the cascade initiated at $n=0$.
@@
Measurement of three non-consecutive shells, such as 6–9–10 or 6–11–12 (Fig. 7), marginally improves the assimilation of faster scales, even beyond the dissipative range, without adverse effects due to gaps. However, excluding a large- or mesoscale shell severely destabilizes the EnKF scheme: configurations measuring only fast scales (e.g., shells 11–12 or 13–14) exhibit blow-up in over $90\%$ of cases. This instability cannot be suppressed by inflation alone and would require prohibitive ensemble sizes or more sophisticated techniques, which are beyond the scope of this work.
@@
This instability can be mitigated by also measuring a large- or mesoscale shell: for instance, adding shell 6 reduces the blow-up rate from $\sim 90\%$ to $\sim 13\%$, with further stabilization achieved by appropriately tuning the inflation factor $\lambda$ (see Appendix B for explanations). Measuring a large- or mesoscale shell promotes the synchronization of traveling bursts originating from large scales and enhances the overall stability of the DA process, consistently with theoretical results for dissipative systems, where a finite number of determining modes suffices to control long-term dynamics [36, 37], and, for example, in the Navier–Stokes context, measuring continuously up to $k=0.2k_{\eta}$ (where $k_{\eta}$ is the dissipative wavenumber) enables the recovery of small-scale dynamics independently of the DA method used [38, 39, 40].
All the results presented in this section are summarized in Table 2 by summing the errors from the first scale (excluding the forced one) up to the dissipative cutoff at scale 15: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$, and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar procedure is applied to the error bars to obtain the total uncertainties $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$, consistently following standard rules for the propagation of experimental errors. In addition, detailed information is provided on the percentage of experiments (out of $N_{\text{exp}}=16$) that required the appropriate adjustment of the inflation strength factor $\lambda$ (see Appendix B).
@@
6, 14 $26.83\pm 1.21$ $28.40\pm 3.23$ $0\%$
4, 14 $25.81\pm 1.88$ $28.74\pm 4.26$ $0\%$
2, 14 $23.15\pm 1.36$ $27.71\pm 4.01$ $0\%$
6 $28.65\pm 1.05$ $27.74\pm 2.47$ $0\%$
6, 7 $7.94\pm 0.43$ $7.29\pm 0.95$ $50\%$
6, 7, 8 $4.85\pm 0.26$ $5.11\pm 0.85$ $6\%$
6, 9, 10 $0.18\pm 0.04$ $0.73\pm 0.53$ $13\%$
6, 11, 12 $0.016\pm 0.004$ $0.04\pm 0.03$ $19\%$
Table 2: The total error is computed by summing the normalized errors from the first scale $n=1$ (excluding the forced one) up to the dissipative cutoff at scale $n=15$: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar summation is applied to the error bars to obtain $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Details on the percentage of experiments (out of $N_{\text{exp}}=16$) that required proper tuning of the inflation strength factor are also provided.
Figure 9: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy with Nudging $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T}$, at varying nudging coefficient $\alpha$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
Figure 10: (a) Normalized ground-truth energy $|u_{n}|^{2}k_{n}^{2/3}$, including signal bursts during the interval $12\leq t/\tau_{0}\leq 14$. Red vertical lines indicate the measured shells ($n=6$, $7$, and $8$). (b) Normalized energy of the Nudging estimate $|\tilde{u}_{n}|^{2}k_{n}^{2/3}$.
@@ VI Comparisons with Nudging
$\left(\frac{d}{dt}+\nu k_{n}^{2}\right)\tilde{u}_{n}(t)=G_{n}[\bm{u}]+f_{n}+\alpha_{n}\delta_{nm}\left[T_{n}(\bm{u},t)-\tilde{u}_{n}(t)\right]\,,$ (19)
@@
Given that measurements are not available at every time step, a continuous linear interpolation $T_{n}(\bm{u},t)$ is introduced:
$T_{n}(\bm{u}(t),t)=\frac{(t_{k+1}-t)z_{n}(t_{k})+(t-t_{k})z_{n}(t_{k+1})}{t_{k+1}-t_{k}},$ (20)
[…]
@@
To ensure that the relaxation rate towards the ground truth scales with observation sparseness, we define the coupling coefficient as $\alpha_{n}=\alpha^{\prime}_{n}/\tau$, where $\tau=dt\Delta t_{\textit{obs}}$. In all experiments using $\Delta t_{\textit{obs}}=\tau_{15}$, the assimilation performance was found to depend only weakly on the specific value of $\alpha^{\prime}_{n}$ on each scale. Therefore, we set $\alpha^{\prime}_{n}=\alpha^{\prime}$ uniformly across all scales. Once this is decided, the way we optimize $\alpha^{\prime}$ depends on the specific set of measured shells.
Figure 11: Comparison between Nudging and EnKF. (a) The structure functions for the labelled cases, for ground-truth $\langle|u_{n}|^{2}\rangle_{T}$, nudged solution $\langle|\tilde{u}_{n}|^{2}\rangle_{T}$ and EnKF $\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}=\tau_{15}=0.002\tau_{0}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(15)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
Figure 10(a) illustrates the effect of varying $\alpha^{\prime}$ on the estimate of the nudging-based energy spectrum estimate $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T}$, compared with the ground truth spectrum $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$, in the case where the shells $n=6,7,8$ are measured. It can be seen that the extreme locality of nudging corrections leads to energy accumulation just before the measured scales, with the magnitude of the effect increasing with $\alpha^{\prime}$. This behavior also reflects in the degradation of the instantaneous reconstruction shown in Fig. 10(b). Combining these two results, we observe that for small $\alpha^{\prime}$ values the energy cascade remains undisturbed, but no effective assimilation occurs. In contrast, high values of $\alpha^{\prime}$ improve reconstruction but lead to accumulation of unphysical energy in the spectrum (see Fig. 10(b)). As a compromise, we identify $\alpha^{\prime}=0.1$ as the ”optimal” value. We also report results for the experiment where scales $n=0,1,2$ are measured, still using $\Delta t_{\textit{obs}}=\tau_{15}$. Consistent with previous discussion, when large scales (including forced shell $n=0$) are observed continuously, energy accumulation is avoided because of immediate synchronization of large-scale bursts. In this case, the optimal choice of $\alpha^{\prime}$ follows the criterion of minimizing the total MSE on the measured shells, $\sum_{n=0}^{2}\langle|u_{n}-\tilde{u}_{n}|^{2}\rangle_{T}$, again yielding $\alpha^{\prime}=0.1$.
@@
Finally, although no significant cascade perturbations are observed when $n=0,1,2$ is measured, Figure 11 confirms that the localized nature of nudging corrections still limits their effectiveness in reconstructing the cascade deeper in the inertial range. In contrast, the EnKF successfully infers up to five additional shells beyond those observed, reaching the independence threshold at $n=10$, compared to $n=5$ for Nudging. In terms of scale separation, this corresponds to a factor of $322^{5}=$ 32.
@@ VII Comparison with En4D-Var
In 4D-Var, assimilation is carried out over a finite time window, within which the observations are distributed, and the forecast trajectory is iteratively adjusted to them by minimizing a cost function. Unlike EnKF, where corrections are applied only at the observation times, in 4D-Var the correction spans the entire assimilation window, making the posterior estimate $\tilde{\bm{u}}^{(i)}(t)$ dependent not only on the observations but also on the window length $T_{a}$ and the iteration step $i$ within the optimization process.
The 4D-Var technique is based on the minimization of a cost function, which in our case is defined as:
$J^{(i)}$ $=$ $\left\|\hat{\bm{u}}(0)-\tilde{\bm{u}}^{(i)}(0)\right\|_{\mathcal{B}}^{2}$ (21)
$+$ $\sum_{t=0}^{T_{a}}\left\|\bm{z}(t)-\bm{H}_{\mathbb{C}}\tilde{\bm{u}}^{(i)}(t)-\bm{v}(t)\right\|_{\bm{R}_{\mathbb{C}}}^{2}\,,$
@@
The variable $\hat{\bm{u}}(0)$ represents the prior (or background) state at the beginning of the window, and $\tilde{\bm{u}}^{(i)}(0)$ is the posterior optimized state (also known as the smoother) obtained after $i$ iterations. In the strong-constraint (i.e. perfect model framework) 4D-Var approach used here, the posterior trajectory is fully determined by its initial condition, i.e., $\tilde{\bm{u}}(t)=\Phi_{0\to t}(\tilde{\bm{u}}(0))$, where $\Phi_{0\to t}$ is the time-marching operator associated with the nonlinear shell model governing equation. The matrix $\mathcal{B}$ is the complex-valued background error covariance, assumed to be given, and $\bm{H}_{\mathbb{C}}$ is the complex extension of the observation operator $\bm{H}$ (see Sec. III). Finally, $\bm{v}(t)$ is a realization of complex zero-mean Gaussian noise with covariance $\bm{R}_{\mathbb{C}}$ (complex version of the matrix in Eq. (13)).
The cost function comprises two terms: the first penalizes the deviation of the posterior from the prior at the initial time, while the second accumulates the innovations, i.e., the discrepancies between the posterior trajectory and the observations over the assimilation window. Solving this optimization directly is computationally intractable even for moderate problem sizes. Therefore, a tangent linear approximation (TLA) is used, where $\Phi_{0\to t}$ is linearized around $(\tilde{\bm{u}}(0))$ and is considered constant within each iteration. Overall, 4D-Var integrates dynamical information over time, resulting in smooth and temporally consistent reconstructions, albeit at the cost of increased computational complexity of iterative nonlinear optimization.
@@
Figure 12: Comparison between En4D-Var (both smoother and filter) and EnKF. (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy with En4D-Var and EnKF $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(15)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.

### v2

@@ V.2 Measured Scale Schemes
In this work we perform a systematic investigation of how the choice of measured scales affects both statistical and instantaneous DA reconstruction, as well as cascading energy transfer. After establishing a synchronization threshold based on observation sparseness $\Delta t_{\textit{obs}}$, we fix it to $\tau_{15}$ (the turnover time of the dissipative scale), which so far has ensured the best assimilation performance, and vary the set of measured shells to assess its impact.
@@
Figure 7: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ (black dots beneath the colored points) and estimated energy $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$ for different sets of measured shells; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(16)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(17)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
@@
A central finding is that measuring two adjacent shells within the inertial range, such as shells 6 and 7, suffices to achieve full synchronization of the large-scale dynamics (Fig. 7(b)). Adding a third measurement (e.g., shells 6, 7, and 8) only assimilates an additional fast scale without significantly enhancing synchronization. In contrast, measuring a single shell (e.g., shell $6$) fails to synchronize the inertial range and disrupts the energy cascade. This disruption arises from the structure of the energy flux, which is mediated through triadic couplings involving terms such as $u_{n}^{*}u_{n+1}^{*}u_{n+2}$ and $u_{n-1}^{*}u_{n}^{*}u_{n+1}$. An inadequate reconstruction of these nonlinear interactions leads to bottlenecks, blocking the transfer of energy from the forced scales and causing its accumulation in the shells immediately preceding the measured one (see Fig. 7(a)). Contiguity between measured shells is therefore critical: adjacent measurements fully constrain the local flux, ensuring seamless energy transfer across scales.
@@
Measurement of three non-consecutive shells, such as 6–9–10 or 6–11–12 (Fig. 7(b) and (c)), substantially improves the assimilation of faster scales (see Table 2), even beyond the dissipative range, without adverse effects due to gaps. However, excluding a large- or mesoscale shell severely destabilizes the EnKF scheme: configurations measuring only fast scales (e.g., shells 11–12 or 13–14) exhibit blow-up in over $90\%$ of cases. This instability cannot be suppressed by inflation alone and would require prohibitive ensemble sizes or more sophisticated techniques, which are beyond the scope of this work.
@@
This instability can be mitigated by also measuring a large- or mesoscale shell: for instance, adding shell 6 reduces the blow-up rate from $\sim 90\%$ to $\sim 13\%$, with further stabilization achieved by appropriately tuning the inflation factor $\lambda$ (see Subsection III.1 for explanations). Measuring a large- or mesoscale shell promotes the synchronization of traveling bursts originating from large scales and enhances the overall stability of the DA process, consistently with theoretical results for dissipative systems, where a finite number of determining modes suffices to control long-term dynamics [43, 44], and, for example, in the Navier–Stokes context, measuring continuously up to $k=0.2k_{\eta}$ (where $k_{\eta}$ is the dissipative wavenumber) enables the recovery of small-scale dynamics independently of the DA method used [45, 46, 47].
All the results presented in this section are summarized in Table 2 by summing the errors from the first scale (excluding the forced one) up to the dissipative cutoff at scale 15: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$, and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar procedure is applied to the error bars to obtain the total uncertainties $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$, consistently following standard rules for the propagation of experimental errors. In addition, detailed information is provided on the number of experiments (out of $N_{\text{exp}}=16$) that required the appropriate adjustment of the inflation strength factor $\lambda$ (see Subsection III.1).
@@
6 $28.65\pm 1.05$ $27.74\pm 2.47$ 0
6, 7 $7.94\pm 0.43$ $7.29\pm 0.95$ 88
6, 7, 8 $4.85\pm 0.26$ $5.11\pm 0.85$ 11
6, 9, 10 $0.18\pm 0.04$ $0.73\pm 0.53$ 22
6, 11, 12 $016\pm 0.004 0.0$$04\pm 0.03$ 33
Table 2: The total error is computed by summing the normalized errors from the first scale $n=1$ (excluding the forced one) up to the dissipative cutoff at scale $n=15$: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar summation is applied to the error bars to obtain $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Details on the number of experiments (out of $N_{\text{exp}}=16$) that required proper tuning of the inflation strength factor are also provided.
Figure 8: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy with Nudging $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T}$, at varying nudging coefficient $\alpha=\alpha^{\prime}/\tau$, with $\tau=\Delta t_{\textit{obs}}=\tau_{15}$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(16)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
Figure 9: (a) Normalized ground-truth energy $|u_{n}|^{2}k_{n}^{2/3}$, including signal bursts during the interval $12\leq t/\tau_{0}\leq 14$. Red vertical lines indicate the measured shells ($n=6$, 77, and 88). (b) Normalized energy of the Nudging estimate $|\tilde{u}_{n}|^{2}k_{n}^{2/3}$.
@@ VI Comparisons with Nudging
$\left(\frac{d}{dt}+\nu k_{n}^{2}\right)\tilde{u}_{n}(t)=G_{n}[\bm{u}]+f_{n}+\alpha_{n}\delta_{nm}\left[T_{n}(\bm{z},t)-\tilde{u}_{n}(t)\right]\,$, (21)
@@
Given that measurements are not available at every time step, a continuous linear interpolation $T_{n}(\bm{z},t)$ is introduced:
$T_{n}(\bm{z},t)=\frac{(t_{k+1}-t)z_{n}(t_{k})+(t-t_{k})z_{n}(t_{k+1})}{t_{k+1}-t_{k}},$ (22)
where $z_{n}(t_{k})=u_{n}(t_{k})+\epsilon_{n}(t_{k})$ and $z_{n}(t_{k+1})=u_{n}(t_{k+1})+\epsilon_{n}(t_{k+1})$ are the ground-truth measurements at times $t_{k}$ and $t_{k+1}$, respectively. The terms $\epsilon_{n}(t_{k})$ and $\epsilon_{n}(t_{k+1})$ are independent measurement errors drawn from a complex zero-mean Gaussian distribution, with covariance $(R_{\mathbb{C}})_{mn}=\delta_{mn}\ 2\ (0.05)^{2}\langle|u_{n}|^{2}\rangle$, where $m,n=0,\ldots,N-1$ index the measured shells (this corresponds to the complex version of the matrix in Eq. (15)).
@@
To ensure that the relaxation rate towards the ground truth scales with observation sparseness, we define the coupling coefficient as $\alpha_{n}=\alpha^{\prime}_{n}/\tau$, where $\tau=\Delta t_{\textit{obs}}$. In all experiments using $\Delta t_{\textit{obs}}=\tau_{15}$, the assimilation performance was found to depend only weakly on the specific value of $\alpha^{\prime}_{n}$ on each scale. Therefore, we set $\alpha^{\prime}_{n}=\alpha^{\prime}$ uniformly across all scales. Once this is decided, the way we optimize $\alpha^{\prime}$ depends on the specific set of measured shells.
Figure 10: Comparison between Nudging and EnKF. (a) The structure functions for the labelled cases, for ground-truth $\langle|u_{n}|^{2}\rangle_{T}$ (black dots beneath the colored points), nudged solution $\langle|\tilde{u}_{n}|^{2}\rangle_{T}$ and EnKF $\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}=\tau_{15}=0.002\tau_{0}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(16)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(17)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
Figure 9(a) illustrates the effect of varying $\alpha^{\prime}$ on the estimate of the nudging-based energy spectrum estimate $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T}$, compared with the ground truth spectrum $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$, in the case where the shells $n=6,7,8$ are measured. It can be seen that the extreme locality of nudging corrections leads to energy accumulation just before the measured scales, with the magnitude of the effect increasing with $\alpha^{\prime}$. This behavior also reflects in the degradation of the instantaneous reconstruction shown in Fig. 9(b). Combining these two results, we observe that for small $\alpha^{\prime}$ values the energy cascade remains undisturbed, but no effective assimilation occurs. In contrast, high values of $\alpha^{\prime}$ improve reconstruction but lead to accumulation of unphysical energy in the spectrum (see Fig. 9(b)). As a compromise, we identify $\alpha^{\prime}=0.1$ as the ”optimal” value.
@@
We also report results for the experiment where scales $n=0,1,2$ are measured. Consistent with previous discussion, when large scales (including forced shell $n=0$) are observed continuously, energy accumulation is avoided because of immediate synchronization of large-scale bursts. In this case, the optimal choice of $\alpha^{\prime}$ follows the criterion of minimizing the total MSE on the measured shells, $\sum_{n=0}^{2}\langle|u_{n}-\tilde{u}_{n}|^{2}\rangle_{T}$, again yielding $\alpha^{\prime}=0.1$. Finally, although no significant cascade perturbations are observed when $n=0,1,2$ is measured, Figure 10 confirms that the localized nature of nudging corrections still limits their effectiveness in reconstructing the cascade deeper in the inertial range. In contrast, the EnKF successfully infers up to five additional shells beyond those observed, reaching the independence threshold at $n=10$, compared to $n=5$ for Nudging. In terms of scale separation, this corresponds to a factor of $2^{5}=32$.
@@ VII Comparison with En4D-Var
In 4D-Var, assimilation is carried out over a finite time window, within which the observations are distributed, and the forecast trajectory is iteratively adjusted to them by minimizing a cost function. Unlike EnKF, where corrections are applied only at the observation times, the 4D-Var is a smoother that adjusts the assimilating model to observations available over an entire assimilation window, making the posterior estimate $\tilde{\bm{u}}(t)$ dependent not only on the observations but also on the window length $T_{a}$.
The posterior estimate $\tilde{\bm{u}}(t)$ in 4D-Var is obtained such that it minimize a cost function defined as:
$J$ $=$ $\left\|\hat{\bm{u}}(0)-\tilde{\bm{u}}(0)\right\|_{\mathcal{B}}^{2}$ (23)
$+$ $\sum_{t=0}^{T_{a}}\left\|\bm{z}(t)-\bm{H}_{\mathbb{C}}\tilde{\bm{u}}(t)-\bm{v}(t)\right\|_{\bm{R}_{\mathbb{C}}}^{2}\,$,
@@
The variable $\hat{\bm{u}}(0)$ represents the prior (or background) state at the beginning of the window, and $\tilde{\bm{u}}(0)$ is the posterior optimized state (also known as the smoother). In the strong-constraint (i.e. perfect model framework) 4D-Var approach used here, the posterior trajectory is fully determined by its initial condition, i.e., $\tilde{\bm{u}}(t)=\Phi_{0\to t}(\tilde{\bm{u}}(0))$, where $\Phi_{0\to t}$ is the time-marching operator associated with the nonlinear shell model governing equation. The matrix $\mathcal{B}$ is the complex-valued background error covariance, assumed to be given, and $\bm{H}_{\mathbb{C}}$ is the complex extension of the observation operator $\bm{H}$ (see Sec. III), needed to produce the complex measurements $z\in\mathbb{C}^{M}$. Finally, $\bm{v}(t)$ is a realization of complex zero-mean Gaussian noise with covariance $\bm{R}_{\mathbb{C}}$ (complex version of the matrix in Eq. (15)).
The cost function comprises two terms: the first penalizes the deviation of the posterior from the prior at the initial time, while the second accumulates the innovations, i.e., the discrepancies between the posterior trajectory and the observations over the assimilation window. Here, we adopt an incremental approach to 4D-Var in which we simplify the cost function by linearizing $\Phi_{0\to t}$ around an initial guess for $\tilde{\bm{u}}(0)$ [22]. This results in a quadratic approximation of the cost function, which is then minimized using Newton’s method. For implementation, we save the linearised operators $\Phi_{0\to t}$ during the forward time-marching, then use them to form the adjoint model, required for obtaining the Jacobian of the cost-function, and explicitly calculate the Hessian, required by the Newton’s method. Once a solution to the simplified cost function is found, the process is repeated by re-linearizing $\Phi_{0\to t}$ around the updated solution. This iterative process continues until convergence is achieved.
@@
Figure 11: Comparison between En4D-Var (both smoother and filter) and EnKF. (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ (black dots beneath the colored points) and estimated energy with En4D-Var and EnKF $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$; the vertical solid line indicates the shell turnover time associated with the assimilation window $\Delta t_{\textit{obs}}$. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(16)) normalized with energies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(17)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.